# MoM: Linear Sequence Modeling with Mixture-of-Memories

### v2

@@ 3.2.2 Linear Recurrent Memory Module
340M Params 15B Tokens L=24, d=1024 Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70 28.81
@@
3.3 Optimizations for MoM
$\bm{k}^{s}_{t}=\bm{x}_{t}\bm{W}^{s}_{k},\bm{v}^{s}_{t}=\bm{x}_{t}\bm{W}^{s}_{v}\in\mathbb{R}^{d}$ (11)
$\bm{M}^{s}_{t}=\bm{M}^{s}_{t-1}+(\bm{k}^{s}_{t})^{T}\bm{v}^{s}_{t}\in\mathbb{R}^{d\times d}$ (12)
$\tilde{\bm{M}}_{t}=\bm{M}^{s}_{t}+\sum_{m}g_{t}^{(m)}\bm{M}^{m}_{t}\in\mathbb{R}^{d\times d}$ (13)
Hardware-efficient Implementation.   Chunk-wise parallel form of linear attention is a computational optimization strategy that divides the input sequence into smaller chunks to enable partial parallel and recursive computation during model training Hua et al. (2022); Yang et al. (2023), which is well supported in open-source frameworks.
In the implementation of MoM, mixing the memories before multiplying by the query is mathematically equivalent to firstly multiplying each memory by the query and then mixing the results. This property allows us to reuse the efficient Triton-based operator implementation of all previous linear sequence modeling methods, facilitate the hardware-efficient training as well as inference of MoM.
Models marked with an asterisk use open-source pretrained weights with identical training configurations.
Scale Model
Wiki.
$\downarrow$
Lamb.
$\downarrow$
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
340M Params 15B Tokens L=24, d=1024 Transformer++ 26.88 76.46 44.91 25.94 34.95 26.90 64.31 51.07 41.35
RetNet 31.07 87.11 44.49 23.04 33.86 23.93 63.49 52.33 40.19
HGRN2 27.90 77.40 45.24 23.63 35.61 24.74 65.45 54.06 41.46
GLA 28.78 79.95 44.53 22.27 34.84 24.94 63.93 51.38 40.32
GSA 28.17 82.50 45.50 24.23 35.00 24.02 64.85 50.43 40.67
Gated DeltaNet 26.47 58.59 46.04 23.55 35.18 27.01 66.05 50.83 41.44
MoM 25.86 55.41 44.65 24.74 36.54 27.93 66.16 51.78 41.97
RetNet 18.18 21.97 57.49 26.88 48.09 37.75 69.37 53.28 48.81
HGRN2 17.32 15.65 58.33 28.07 51.93 42.31 71.33 52.01 50.66
GLA 17.61 19.66 55.18 27.56 48.89 40.03 69.86 53.91 49.24
GSA 16.69 16.02 58.33 28.33 50.98 42.03 72.25 53.43 50.89
Gated DeltaNet 17.14 18.80 56.82 27.39 49.77 39.94 71.76 51.78 49.58
MoM 16.64 14.83 55.35 27.99 50.95 43.43 71.27 56.83 50.97
4 Experiments
4.1 Experimental Setups
Models.   In our experiments, we employ the Gated DeltaNet Yang et al. (2024) as the memory update mechanism in MoM. The model is configured with four memory states, two of which are activated at each time step, along with a shared memory.
Baselines.   We evaluate MoM against several linear recurrent models and Transformers, including RetNet Sun et al. (2023), GLA Yang et al. (2023), Gated DeltaNet Yang et al. (2024), and Transformer++ Touvron et al. (2023), which incorporates Rotary Position Embeddings Su et al. (2024) and GLU Shazeer (2020) into the Transformer architecture. To ensure a fair comparison, we train all baseline models from scratch using the exact same number of tokens.
Training.   We follow the training procedure described by Yang et al. (2023), utilizing the SlimPajama dataset Soboleva et al. (2023) sampled with 100B tokens and tokenized using the Mistral tokenizer Jiang et al. (2023). We train models from scratch with parameter sizes of 340M and 1.3B, respectively. For the 340M models, we train on 15B tokens with a batch size of 0.5M tokens. The warmup tokens count is set to 0.25M. We set the hidden ratio of our model to 3 to keep the activated parameter count approximately the same. For the 1.3B models, we train on 100B tokens with a batch size of 2M tokens. We utilized publicly available pretrained weights from Zhang et al. (2024) with exactly same configuration. The warmup tokens count is 1B. We employ AdamW optimizer Loshchilov et al. (2017); Sun et al. (2024b) with learning rate of 3e-4 with cosine learning rate schedule Zhou et al. (2020). The weight decay is set to 0.01 and gradient clipping is 1.0.
Parameter Explanation. The parameter counts differ across models due to the varying modules they contain. For reference, the "340M" model refers to a configuration with 24 layers, each having a hidden size of 1024. Similarly, "1.3B" refers to 24 layers with a hidden size of 2048. Actual parameter variations may occur due to differences in specific linear modules. The primary goal of MoM is to enable sparse activation and expand the memory capacity of linear sequence models. To achieve this, we applied sparse activation only to the key and value projections, resulting in a minor increase in activated parameters. This small increase is well justified given the substantial improvements in performance. We’ve also conducted comparisons with models that traditionally use larger parameter counts to expand memory capacity. A detailed discussion on fairness is provided in Appendix F.
4.2 Main Results
@@ 4.2.1 Recall-intensive Tasks
To thoroughly assess our model’s proficiency in such scenarios, we test six recall-intensive tasks following Arora et al. (2024): FDA Arora et al. (2023), SWDE Arora et al. (2023); Lockard et al. (2019), SQuAD Rajpurkar et al. (2018), NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017) and Drop Dua et al. (2019). These tasks are designed to challenge a model’s ability to perform context-based retrieval and comprehension.
As shown in Table 3.2.2, our proposed approach, benefiting from increased memory capacity and memory mixing mechanism, achieves significant improvements over other linear sequence models. Specifically, our model effectively narrows the performance gap with Transformer models. This improvement underscores the advantage of our method in capturing and utilizing long-range dependencies, thereby enhancing performance on tasks that require extensive contextual understanding.
4.2.2 Commonsense Reasoning Tasks
As shown in Table 3.3, we report the language modeling perplexity and zero-shot performance of commonsense reasoning tasks following Zhang et al. (2024) which includes WikiText Merity et al. (2016), LAMBADA Paperno et al. (2016), ARC-easy, ARC-challenge Clark et al. (2018), HellaSwag Zellers et al. (2019), PiQA Bisk et al. (2020) and WinoGrande Sakaguchi et al. (2019). The evaluation results are based on the lm-evaluation-harness Gao et al. (2024).
Model Sum FS Syn Code Avg.
RetNet 6.30 15.76 2.64 40.52 13.61
HGRN2 6.51 15.50 2.61 40.11 13.02
GSA 7.75 20.29 1.92 42.83 14.61
Gated DeltaNet 7.14 18.00 2.10 41.52 13.98
MoM 6.89 21.26 2.63 47.79 15.64
Table 4: LongBench Benchmark Results. All evaluations were done using OpenCompass. Note: Sum = Summarization, FS = Few-shot, Syn = Synthetic.
Model FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70
MoM 22.98 29.90 29.69 16.60 48.82 20.99 28.16
MoM Hybrid 58.13 44.05 35.71 20.18 48.10 20.60 37.80
Model
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
Transformer++ 44.91 25.94 34.95 26.90 64.31 51.07 41.35
MoM 44.65 24.74 36.54 27.93 66.16 51.78 41.97
MoM Hybrid 46.55 24.49 36.45 28.86 65.51 52.41 42.38
@@ 4.2.3 Long Context Tasks
Assessing performance on long-context tasks is crucial for linear models, as it reflects their ability to handle long-range dependencies effectively. We evaluated our model’s comprehension of long contexts using the LongBench benchmark Bai et al. (2024); Contributors (2023). In Table 4, we present the average results across various categories, including summarization, few-shot learning, synthetic tasks, and code completion, along with the overall mean across all tasks. The complete detailed results are provided in Appendix 10.
@@ 4.2.4 The Hybrid of MoM and Transformer
To validate the effectiveness of our mixed memory mechanism, we compare our MoM model with mixed memories to a baseline model that uses an expanded single memory with the same activated memory capacity. We adopt the same memory update method as existing linear models and extend it within our MoM framework. For comparison, we employed the commonly used method of expanding the single memory by expanding the dimension of $v$ to match the total size of all activated memories in the MoM model. Note that while our MoM model activates multiple memories during retrieval, the mixing process effectively consolidates these into a memory capacity equivalent to a single memory. In contrast, the expanded single memory approach increases the entire retrieval capacity which leads to a significant increase in the number of parameters. We evaluate their performance on common-sense reasoning tasks and recall-intensive tasks in Table 6.
The experimental results demonstrated that using multiple mixed memories leads to a greater improvement than simply expanding the capacity of a single memory. This confirms that mixed memory can effectively reduce interference from different inputs. Assigning inputs specifically to different memories, combined with the use of a forget gate, proves to be a more effective approach for reducing interference than relying solely on a forget gate.
Figure 2: Inference Efficiency of MoM. We demonstrate the inference (decoding) time and GPU memory consumption required to generate 1K tokens at specific sequence lengths.
@@ 4.2.7 Length Extrapolation
We pretrained the models on the Slimpajama dataset with a 2K context length and conducted extrapolation experiments on various lengths using the FinewebPenedo et al. (2024) dataset. We extended the length to 32K to calculate perplexity (ppl). As shown in Fig 3, the Transformer model experienced a significant increase in ppl due to its poor extrapolation capability. Among the linear models, MoM achieved the best results.
Figure 3: Length Extrapolation. We extrapolated models trained on 2K sequences to a length of 32K for perplexity (ppl) evaluation. The ppl of Transformer++ rises sharply from 13.4 at 2k to beyond the display range, reaching 2271.3 at 32k.
Figure 4: Memory Load Balance Analysis. Token Routing Distribution Across Layers and Memories with and without Aux Loss.
@@ 4.2.8 Memory Analysis
Memory Load Balance Analysis. To evaluate whether each memory segment in MoM is effectively balanced during inference on downstream tasks, we analyzed the number of tokens routed to each layer using around 300k tokens from the ARC-easy benchmark. We visualized the results with and without auxiliary loss in Fig 4 with heatmaps. Due to the adoption of auxiliary loss, the memory segments in each layer are almost uniformly routed and activated.
Functions of Different Memories. We analyzed the memory routing within the model and identified the categories of tokens processed by each of the four memories. By examining an intermediate layer, we observed that, unlike traditional MoE in FFN layers where experts often lack clear specialization and focus on syntactic punctuationJiang et al. (2023), MoM displays distinct specialization among its memories. This finding suggests potential for exploring an MoE architecture where each memory serves a specific role. These distinctions among different memories are presented in Table 7.
4.2.9 Ablation Study
To evaluate the influence of memory hyperparameters in our MoM model, we conducted ablation studies using a 340 million parameter model trained on 15 billion tokens. Our investigation primarily focused on the impact of the number of memories, the number of activations, and the use of shared memory on model performance. The results of these experiments are presented in Table 8.
@@ 4 Interrogative words/incomplete nouns Fragmented/open information
Acc $\uparrow$
Number of Memories and Activations
(all cases include shared memory by default)
Memories Activated
2 1
25.05
3 2
27.31
4 2
27.59
8 2
25.56
@@
w/ shared memory 27.59
w/o shared memory 26.06
Table 8: Ablation Study. We performed ablation studies on the number of memories and the use of shared memory. The table presents the average results across all recall-intensive tasks.
@@ 5 Conclusion
In this paper, we propose Mixture-of-Memories (MoM), a novel architecture that enhances memory capacity and eliminates memory interference, inspired by the brain’s mechanisms for long-term memory retention. By leveraging multiple independent memory states, MoM significantly improves performance on recall-intensive tasks while maintaining the efficiency advantages of linear-time training and constant-memory inference. Instead of simply discarding tokens as done in gating mechanisms, our memory separation paradigm provides a more effective way to preserve sequence information. Our experimental results demonstrate that MoM outperforms existing linear sequence modeling methods, particularly on tasks requiring strong recall, and achieves performance comparable to Transformer models. This makes MoM a promising approach for applications need strong efficiency and recall-intensive performance, paving the way for sparse solutions in sequence modeling.
6 Acknowledgments
This work is supported by the Shanghai Artificial Intelligence Laboratory.
@@ Appendix A Related Work
Linear recurrent models, comprising linear attention, linear RNNs, state-space models (SSMs), have garnered significant research interests Qin et al. (2023b). The advancement of SSMs began with the pioneering work on S4 Gu et al. (2022), which was later optimized through a diagonalized version Gupta et al. (2022). Despite their strong performance on the LRA benchmark, these models have faced challenges in language modeling mainly because they rely solely on data-independent processes. As research progressed, constant forgetting gates were introduced, helping to alleviate some interference issues by uniformly managing memory decay Sun et al. (2023); Gu and Dao (2024). The next breakthrough involved data-dependent forget gates. These allowed models to dynamically adjust memory updates based on the input data, significantly enhancing performance across various tasks Qin et al. (2024c, d); Yang et al. (2023); Zhang et al. (2024); Yang et al. (2024); Qin et al. (2024a). Sequence parallelism techniques are well adapted on linear recurrent models Sun et al. (2024a, 2025) for efficient long context training. There have also been recent advancements in scaling law and test-time regression optimization Shen et al. (2024); Sun et al. (2024c); Behrouz et al. (2024).
@@
Mixture-of-Experts (MoE) is a technique designed to enhance the capacity of deep neural networks while maintaining computational efficiency Fedus et al. (2022); Rajbhandari et al. (2020); Lepikhin et al. (2020); Tang et al. (2023); Zhu et al. (2024); Qu et al. (2024). MoE achieves this by activating a subset of parameters, known as "experts", for each input, which reduces the computational costs. Shazeer first integrated MoE into LSTM layers Shazeer et al. (2017). The Switch Transformer Fedus et al. (2022) refined this approach by simplifying the gating mechanism to select only one expert per input. Gshard Lepikhin et al. (2020) further advanced this by using a top-2 expert routing strategy to improve performance. Recent MoE models, such as Deepseek-MoE Dai et al. (2024), introduce shared experts to capture and consolidate common knowledge across different contexts, while designing fine-grained experts to increase combinatorial flexibility.
@@ Appendix B Comparison between MoM and MoE
  • Structure: In conventional MoE, each expert is a separate neural network within the feedforward network (FFN) layer Team (2024). In contrast, in MoM, each memory is an RNN state with unique key-value projection weights to generate different key-value pairs. MoE operates during the channel mixing phase, where each token is processed independently by selected experts. On the other hand, MoM functions during the token mixing phase, where each memory processes different segments of the sequence, preserving inter-token relationships.
Appendix C Dataset and Benchmark
We pretrained model on SlimPajama dataset. For 340M models, we train on a sample of 15B tokens. For 1.3B models, we train on a sample of 100B tokens.
  • SlimPajama Soboleva et al. (2023) is a high-quality, optimized subset of the RedPajama dataset, designed for large-scale language model training. It includes diverse text sources such as Common Crawl, Wikipedia, books, and GitHub code, with a primary focus on English. The dataset is cleaned, deduplicated, and optimized for efficiency and performance.
For the benchmark, we tested on these tasks:
  • WikiText Merity et al. (2016): A dataset consisting of high-quality Wikipedia articles, primarily in English, designed for language modeling tasks with 62 test samples. The text covers a wide range of topics, including history, science, and culture, and is authored by Wikipedia contributors, who come from diverse demographic backgrounds.
  • LAMBADA Paperno et al. (2016): An English-language dataset for evaluating contextual understanding in language models with 5153 test samples. It consists of narrative texts sourced from books, requiring models to predict the final word of a passage. The data reflects a mix of literary styles and author demographics.
  • ARC-Easy & ARC-Challenge Clark et al. (2018): A set of multiple-choice science questions in English, sourced from standardized exams and educational materials with 2376 and 1172 test samples. The dataset represents the domain of elementary and high school science, with questions authored by educators and test designers. ARC-Easy includes straightforward questions, while ARC-Challenge contains more difficult ones that require advanced reasoning.
  • HellaSwag Zellers et al. (2019): An English-language dataset designed for commonsense reasoning, where models must choose the most plausible continuation of a sentence. The text is derived from activity descriptions (e.g., WikiHow), covering everyday scenarios. The dataset was constructed adversarially to be challenging for language models. It has 10003 test samples.
  • PiQA Bisk et al. (2020): A dataset focused on physical commonsense reasoning in English with 3084 test samples. The text consists of everyday tasks and scenarios, requiring models to determine the most practical way to perform an action. The data is sourced from crowdsourced descriptions, reflecting a broad range of common human experiences.
  • WinoGrande Sakaguchi et al. (2019): A large-scale English dataset for commonsense reasoning, based on the Winograd Schema Challenge with 1267 test samples. It tests pronoun resolution in ambiguous contexts, with sentences sourced and refined through crowdsourcing. The dataset aims to reduce annotation biases by diversifying sentence structures and topics.
  • FDA Arora et al. (2024, 2023): An English-language dataset includes 100 randomly sampled PDF files of FDA 510(k) pre-market notification submissions, each with 16 manually annotated attributes like device classification and predicate device codes. It has 1102 test samples.
  • SWDE Arora et al. (2024, 2023); Lockard et al. (2019): An English dataset dataset is used for evaluating information extraction systems. It comprises data from 8 movie websites (e.g., IMDB, Rotten Tomatoes) and 5 university websites (e.g., US News), with each site containing 1063 to 2000 annotated web pages and $1$ test samples.
  • SQuAD Rajpurkar et al. (2018): A reading comprehension dataset with English questions based on Wikipedia articles with 2984 test samples. Answers are text segments from the articles or the questions can be unanswerable. SQuAD2.0 adds over 50,000 unanswerable questions to the original 100,000 from SQuAD1.1, requiring systems to both answer questions and identify when no answer is supported.
  • NQ Kwiatkowski et al. (2019): A large-scale English dataset for open-domain question answering, where questions are sourced from real Google search queries. Answers are extracted from Wikipedia articles, reflecting general knowledge across various domains. It has 3157 test samples
  • TriviaQA Joshi et al. (2017): An English-language question-answering dataset consisting of trivia-style questions with 1688 test samples. The dataset includes both human-generated questions and web-sourced evidence, covering a broad range of topics such as history, science, and entertainment.
  • DROP Dua et al. (2019): A reading comprehension benchmark in English that requires discrete reasoning, such as arithmetic and logical operations, over passages. The dataset is sourced from Wikipedia and challenges models to go beyond simple span extraction by performing multi-step reasoning. It has 2087 test samples.
  • LongBench Bai et al. (2024) is a bilingual (Chinese and English) benchmark for evaluating long-context understanding in large language models. It covers 21 tasks across six categories, including QA, summarization, and code completion with 4750 test samples.
All datasets used in this work are publicly available and have been released by their original creators, who are responsible for ensuring privacy protection. These datasets are used in accordance with their respective licenses and intended purposes. No modifications or derivative datasets have been created.
Appendix D Experiments Details
Our experiments were conducted using 32 NVIDIA A800 GPUs. Training the 340M parameter model required approximately 10 hours, while the 1.3B parameter model took around 6 days. All models were trained and evaluated using a fixed random seed of 42 to ensure reproducibility. The reported results are based on a single run, without aggregation over multiple random seeds.

### v3

@@ 4.2 Main Results
380M Params 15B Tokens L=24, d=1024 Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70 28.81
@@ 4.2.1 Recall-intensive Tasks
To thoroughly assess our model’s proficiency in such scenarios, we test six recall-intensive tasks following Arora et al. (2024): FDA (Arora et al., 2023), SWDE (Arora et al., 2023; Lockard et al., 2019), SQuAD (Rajpurkar et al., 2018), NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017) and Drop (Dua et al., 2019). These tasks are designed to challenge a model’s ability to perform context-based retrieval and comprehension.
As shown in Table 2, our proposed approach, benefiting from increased memory capacity and memory mixing mechanism, achieves significant improvements over other linear sequence models. Specifically, our model effectively narrows the performance gap with Transformer models. This improvement underscores the advantage of our method in capturing and utilizing long-range dependencies, thereby enhancing performance on tasks that require extensive contextual understanding.
Table 3: LongBench Benchmark Results. Note: Sum = Summarization, FS = Few-shot, Syn = Synthetic.
Model Sum FS Syn Code Avg. RetNet 6.30 15.76 2.64 40.52 13.61 HGRN2 6.51 15.50 2.61 40.11 13.02 GSA 7.75 20.29 1.92 42.83 14.61 Gated DeltaNet 7.14 18.00 2.10 41.52 13.98 MoM 6.89 21.26 2.63 47.79 15.64
@@ 4.2.2 Long Context Tasks
Assessing performance on long-context tasks is crucial for linear models, as it reflects their ability to handle long-range dependencies effectively. We evaluated our model’s comprehension of long contexts using the LongBench benchmark (Bai et al., 2024; Contributors, 2023). In Table 3, we present the average results across various categories, including summarization, few-shot learning, synthetic tasks, and code completion, along with the overall mean across all tasks. The complete detailed results are provided in Appendix I.
@@
To validate the effectiveness of our mixed memory mechanism, we compare our MoM model with mixed memories to a baseline model that uses an expanded single memory with the same activated memory capacity. We adopt the same memory update method as existing linear models and extend it within our MoM framework. For comparison, we employed the commonly used method of expanding the single memory by expanding the dimension of $\bm{v}$ to match the total size of all activated memories in the MoM model. We evaluate their performance on common-sense reasoning tasks and recall-intensive tasks in Table 4.
The experimental results demonstrated that using multiple mixed memories leads to a greater improvement than simply expanding the capacity of a single memory with less parameters. This confirms that mixed memory can effectively reduce interference from different inputs. Assigning inputs specifically to different memories, combined with the use of a forget gate, proves to be a more effective approach for reducing interference than relying solely on a forget gate.
@@ 4.2.5 Length Extrapolation
We pretrained the models on the Slimpajama dataset with a 2K context length and conducted extrapolation experiments on various lengths using the Fineweb(Penedo et al., 2024) dataset. We extended the length to 32K to calculate perplexity (ppl). As shown in Fig 4, the Transformer model experienced a significant increase in ppl due to its poor extrapolation capability. Among the linear models, MoM achieved the best results.
[Uncaptioned image] Figure 3: Inference Efficiency of MoM. We demonstrate the inference time and GPU memory consumption required to generate 1K tokens at specific sequence lengths. [Uncaptioned image] Figure 4: Length Extrapolation. We extrapolated models trained on 2K sequences to a length of 32K for perplexity (ppl) evaluation.
@@ 4.2.6 Memory Analysis
Memory Load Balance Analysis. To evaluate whether each memory segment in MoM is effectively balanced during inference on downstream tasks, we analyzed the number of tokens routed to each layer using around 300k tokens from the ARC-easy benchmark. We visualized the results with auxiliary loss (following the formulation introduced in Switch Transformer Fedus et al. (2022)) in Fig 5 with heatmaps and we also visualized results with auxiliary loss in Fig 9. Due to the adoption of auxiliary loss, the memory segments in each layer are almost uniformly routed and activated.
Functions of Different Memories. We analyzed the memory routing within the model and identified the categories of tokens processed by each of the four memories. By examining an intermediate layer, we observed that, unlike traditional MoE in FFN layers where experts often lack clear specialization and focus on syntactic punctuation(Jiang et al., 2023), MoM displays specialization among its memories. This finding suggests potential for exploring an MoE architecture where each memory serves a specific role. These distinctions among different memories are presented in Table 5.
@@ 4 Interrogative words/incomplete nouns Fragmented/open information
Figure 5: Memory Load Balance Analysis. Token Routing Distribution Across Layers and Memories with Aux Loss.
4.2.7 MoM Scaling Up & Ablation Study
We examine the effect of scaling both the number of memory states and the number of top-$k$ activations in MoM. To ensure comparability, Fig. 6 reports results with a fixed activation ratio of 0.5. Increasing the number of memories from 1 to 8 consistently improves performance across both recall-intensive and commonsense benchmarks. These results indicate that enlarging the memory pool effectively mitigates interference and enhances capacity. More comprehensive results covering other activation ratios and activation settings are provided in Appendix H.
We further study the influence of auxiliary loss and shared memory in MoM, using a 380M-parameter model trained on 15B tokens. As shown in Table 6, auxiliary loss improves stability and performance when applied with a suitable weight. In addition, shared memory consistently benefits performance with global information. These results highlight the complementary roles of auxiliary loss and shared memory in stabilizing and enhancing MoM.
Figure 6: Scaling performance with increasing number of memories with a fixed activation ratio of 0.5.
Table 6: Ablation on memory count and shared memory, showing average accuracy across recall-intensive tasks.
Recall $\uparrow$ Common $\uparrow$
Aux Loss Scale
1e-2 27.59 42.10
5e-3 26.55 41.71
1e-3 28.16 41.97
0 27.23 41.58
@@ 4.2.7 MoM Scaling Up & Ablation Study
w/ shared memory 28.16 41.97
w/o shared memory 26.06 40.38
@@ 5 Conclusion
In this paper, we propose MoM, a novel architecture that enhances memory capacity and eliminates memory interference. By leveraging multiple independent memory states, MoM significantly improves performance on recall-intensive tasks while maintaining the efficiency advantages of linear models. Instead of simply discarding tokens as done in gating mechanisms, our memory separation paradigm provides a more effective way to preserve sequence information. Our experimental results demonstrate that MoM outperforms existing linear sequence modeling methods, particularly on tasks requiring strong recall, and achieves performance comparable to Transformer models. This makes MoM a promising approach for applications need strong efficiency and recall-intensive performance, paving the way for efficient sequence modeling.
6 Ethics Statement
This work does not involve human subjects, sensitive data, or high-risk applications. All experiments are conducted on publicly available datasets. We encourage responsible and ethical use of the proposed methods in line with community standards.
@@ Appendix A Related Work
Linear recurrent models, comprising linear attention, linear RNNs, state-space models (SSMs), have garnered significant research interests (Qin et al., 2023b). The advancement of SSMs began with the pioneering work on S4 (Gu et al., 2022), which was later optimized through a diagonalized version (Gupta et al., 2022). Despite their strong performance on the LRA benchmark, these models have faced challenges in language modeling mainly because they rely solely on data-independent processes. As research progressed, constant forgetting gates were introduced, helping to alleviate some interference issues by uniformly managing memory decay (Sun et al., 2023; Gu & Dao, 2024). The next breakthrough involved data-dependent forget gates. These allowed models to dynamically adjust memory updates based on the input data, significantly enhancing performance across various tasks (Qin et al., 2024c; d; Yang et al., 2023; Zhang et al., 2024; Yang et al., 2024; Qin et al., 2024a). Sequence parallelism techniques are well adapted on linear recurrent models (Sun et al., 2024a; 2025) for efficient long context training. There have also been recent advancements in scaling law and test-time regression optimization (Shen et al., 2024; Sun et al., 2024c; Behrouz et al., 2024).
@@
Mixture-of-Experts (MoE) is a technique designed to enhance the capacity of deep neural networks while maintaining computational efficiency (Fedus et al., 2022; Rajbhandari et al., 2020; Lepikhin et al., 2020; Tang et al., 2023; Zhu et al., 2024; Qu et al., 2024). MoE achieves this by activating a subset of parameters, known as “experts”, for each input, which reduces the computational costs. Shazeer first integrated MoE into LSTM layers (Shazeer et al., 2017). The Switch Transformer (Fedus et al., 2022) refined this approach by simplifying the gating mechanism to select only one expert per input. Gshard (Lepikhin et al., 2020) further advanced this by using a top-2 expert routing strategy to improve performance. Recent MoE models, such as Deepseek-MoE (Dai et al., 2024), introduce shared experts to capture and consolidate common knowledge across different contexts, while designing fine-grained experts to increase combinatorial flexibility.
@@ Appendix B Comparison between MoM and MoE
  • Structure: In conventional MoE, each expert is a separate neural network within the feedforward network (FFN) layer such as Qwen-MoE (Team, 2024) and Linear-MoE Sun et al. (2024a). In contrast, in MoM, each memory is an RNN state with unique key-value projection weights to generate different key-value pairs. MoE operates during the channel mixing phase, where each token is processed independently by selected experts. On the other hand, MoM functions during the token mixing phase, where each memory processes different segments of the sequence, preserving inter-token relationships.
Appendix C Experiments Details
For the 380M models, we train on 15B tokens with a batch size of 0.5M tokens. The warmup tokens count is set to 0.25M. We set the hidden ratio of our model to 3 to keep the activated parameter count approximately the same. For the 1.3B models, we train on 100B tokens with a batch size of 2M tokens. The warmup tokens count is 1B. We employ AdamW optimizer (Loshchilov et al., 2017; Sun et al., 2024b) with learning rate of 3e-4 with cosine learning rate schedule (Zhou et al., 2020). The weight decay is set to 0.01 and gradient clipping is 1.0. Our experiments were conducted using 32 NVIDIA A800 GPUs. Training the 380M parameter model required approximately 10 hours, while the 1.3B parameter model took around 6 days.
Appendix D Commonsense Reasoning Tasks
As shown in Table 7, we report the language modeling perplexity and zero-shot performance of commonsense reasoning tasks following (Zhang et al., 2024) which includes WikiText (Merity et al., 2016), LAMBADA (Paperno et al., 2016), ARC-easy, ARC-challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PiQA (Bisk et al., 2020) and WinoGrande (Sakaguchi et al., 2019). The evaluation results are based on the lm-evaluation-harness (Gao et al., 2024).
Scale Model
Wiki.
[…]
Lamb.
[…]
ARC-e
[…]
ARC-c
$\mathrm{acc_{n}}\uparrow$
Hella.
$\mathrm{acc_{n}}\uparrow$
Lamb.
$\mathrm{acc}\uparrow$
PIQA
$\mathrm{acc}\uparrow$
Wino.
$\mathrm{acc}\uparrow$
Avg.
380M Params 15B Tokens L=24, d=1024 Transformer++ 26.88 76.46 44.91 25.94 34.95 26.90 64.31 51.07 41.35
RetNet 31.07 87.11 44.49 23.04 33.86 23.93 63.49 52.33 40.19
HGRN2 27.90 77.40 45.24 23.63 35.61 24.74 65.45 54.06 41.46
GLA 28.78 79.95 44.53 22.27 34.84 24.94 63.93 51.38 40.32
GSA 28.17 82.50 45.50 24.23 35.00 24.02 64.85 50.43 40.67
Gated DeltaNet 26.47 58.59 46.04 23.55 35.18 27.01 66.05 50.83 41.44
MoM 25.86 55.41 44.65 24.74 36.54 27.93 66.16 51.78 41.97
RetNet 18.18 21.97 57.49 26.88 48.09 37.75 69.37 53.28 48.81
HGRN2 17.32 15.65 58.33 28.07 51.93 42.31 71.33 52.01 50.66
GLA 17.61 19.66 55.18 27.56 48.89 40.03 69.86 53.91 49.24
GSA 16.69 16.02 58.33 28.33 50.98 42.03 72.25 53.43 50.89
Gated DeltaNet 17.14 18.80 56.82 27.39 49.77 39.94 71.76 51.78 49.58
MoM 16.64 14.83 55.35 27.99 50.95 43.43 71.27 56.83 50.97